My question was 13th root x - 7th root x / 5th root x - cube root x !!!

Grade:Upto college level

1 Answers

Anurag Kishore
13 years ago

Hi

limx->1( x1/13 - x1/7)/( x1/5 - x1/3)                    (0/0 form    so use L hospitals rule)

= (1/13 - 1/7)/(1/5 - 1/3) = 45/91

 

 

 

Thanks

Anurag